# On the Vulnerability of Applying Retrieval-Augmented Generation within Knowledge-Intensive Application Domains

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has been empirically shown to enhance the performance of large language models (LLMs) in knowledge-intensive domains such as healthcare, finance, and legal contexts. Given a query, RAG retrieves relevant documents from a corpus and integrates them into the LLMs’ generation process. In this study, we investigate the adversarial robustness of RAG, focusing specifically on examining the retrieval system. First, across 225 different setup combinations of corpus, retriever, query, and targeted information, we show that retrieval systems are vulnerable to universal poisoning attacks in medical Q&A. In such attacks, adversaries generate poisoned documents containing a broad spectrum of targeted information, such as personally identifiable information. When these poisoned documents are inserted into a corpus, they can be accurately retrieved by any users, as long as attacker-specified queries are used. To understand this vulnerability, we discovered that the deviation from the query’s embedding to that of the poisoned document tends to follow a pattern in which the high similarity between the poisoned document and the query is retained, thereby enabling precise retrieval. Based on these findings, we develop a new detection-based defense to ensure the safe use of RAG. Through extensive experiments spanning various Q&A domains, we observed that our proposed method consistently achieves excellent detection rates in nearly all cases.
@@ 1 Introduction
Large Language Models (LLMs) have achieved exceptional performance across a wide range of benchmark tasks spanning multiple domains [1, 2, 3]. However, they also suffer from several undesirable behaviors. For example, LLMs can generate responses that seem reasonable but are not factually correct, a phenomenon known as hallucination [4]. Additionally, due to data privacy regulations such as GDPR [5], direct training on specific data domains may be restricted. This can result in disparities between their acquired internal knowledge and the real-world challenges they encounter, leading to unreliable generation. These challenges can be particularly concerning in domains require extensive knowledge, such as healthcare [6, 7], finance [8] and legal question-answering [9].
Retrieval-Augmented Generation (RAG) [10, 11, 12, 13, 14] has emerged as a promising solution to these challenges by integrating external knowledge to LLMs’ generations. The RAG approach typically involves two steps: retrieval and augmentation. Upon receiving an input query, RAG retrieves the top $K$ relevant data from an external data corpus. It then integrates this retrieved information with its internal knowledge to make final predictions. Empirical evidence suggests that LLMs employing the RAG scheme significantly outperform their non-retrieval-based counterparts in knowledge-intensive domains like finance and medicine [12, 14]. For instance, the authors of [14] developed a state-of-the-art benchmark for the use of RAG in the medical domain. The authors observed an increase in prediction accuracy of up to 18% with RAG compared to non-retrieval and chain-of-thoughts versions across large-scale healthcare tasks, utilizing 41 different combinations of medical data corpora, retrievers, and LLMs.
@@ 1.1 Main Contributions
Revealing the safety risks for retrieval systems: A case study for medical Q&A. We demonstrate that dense retrieval systems are vulnerable to what we term as ‘universal poisoning attacks’ in medical Q&A, across a total of 225 different use-case combinations of corpus, retriever, query, and targeted information. As shown in Figure 1 below, in these attacks, adversaries can append nearly every sort of information, such as personally identifiable information (PII) and adversarial treatment recommendation, to a set of attacker-specified queries. Once these poisoned documents are injected into a large-scale corpus, such as Wikipedia and PubMed, they can be accurately retrieved, often with high rankings, e.g., top $1$, using attacker-specified queries. Depending on attackers’ goals, these documents will lead to safety risks such as (1) leakage of PII, (2) adversarial recommendations for treatments, and (3) jailbreaking the LLM during the inference stage once they are used as context.
@@
Figure 2: Defense methods against universal poisoning attacks. Clean documents (collected from the Textbook [20] corpus) and poisoned documents are indistinguishable using popular defenses, such as the $\ell_{2}$-norm of their embeddings. However, under our proposed distance measurements, a clear separation between clean and poisoned documents is observed (right-most figure).
@@ 1.2 Related Work
Retrieval-Augmented Generation (RAG) RAG, popularized by [11], is a widely adopted approach that integrates retrieval-based mechanisms into generative models to improve performance across various language tasks [21]. The basic flow of RAG follows a ‘retrieve and read’ fashion [11, 22, 23, 12, 13, 10, 24, 25, 26]. In this fashion, given an input query, one first retrieves relevant data from the external data corpus and then employs generative models as ‘readers’, for example, by directly appending the retrieved documents to the queries as context and then feeding them as a whole to LLMs for making predictions. Approaches for enhancing the efficient use of retrieved documents to improve the LLMs’ reading capability include using chunked cross-attention during generation [12], prompt-tuning [23], and in-context learning [13].
@@
Dense retrieval systems There are two main categories of retrievers: sparse retrieval systems, which match through lexical patterns (e.g., BM25) [27, 28], and deep neural network-based dense retrievers, which match through semantic meanings [29, 30, 31]. Due to the near-exact token-level matching pattern, sparse retrievers’ performance has been shown to be worse than that of dense retrievers in several domains [32], such as healthcare [14]. There are two popular approaches for training dense retrievers: supervised [33, 34] and self-supervised [29, 31, 30]. In the supervised regime, given a paired query and document, the goal is to maximize the similarity, i.e., inner product, between their embeddings. Motivated by advances in unsupervised learning, recent methods have started to apply contrastive learning [35] for training, where they observed improved performance across multiple benchmarks. In our work, in addition to popular retrievers, we also consider a retriever trained with contrastive learning on medical datasets: MedCPT [31].
@@
Adversarial attacks against RAG/retrieval systems Current approaches for attacking RAG [15, 16] mostly involve poisoning the corpus to deceive the retrieval system [18, 36] into retrieving the poisoned documents for adversarial purposes. The recent work of PoisonedRAG aims to backdoor the RAG by tricking the LLM into generating attacker-specified answers based on the retrieved attacker-crafted documents [15]. The authors develop both black-box and white-box attacks to achieve this goal. In black-box methods, they directly append the context for answering the question to the query and inject it into the corpus, similar to our method in implementation. However, their approaches differ from ours in terms of goals, insights, and application scenarios. First, our objective is to investigate and comprehend the robustness of retrieval systems employed in RAG. We achieve this by injecting various types of information, encompassing both irrelevant and relevant content, into the corpus and evaluating the ease or difficulty of retrieval. Their goal, on the other hand, is to inject only query-relevant context to deceive the LLM’s generation process upon retrieval. Second, in terms of insights, we provide explanations on the difficulty/easiness of the retrieval of different kinds of information, which is not covered in their work. Third, we focus on the application domain of healthcare, while they focus on the general Q&A setting.
@@ 2 Preliminary and Threat Model
Notations Denote the set of vocabulary to be considered as $\mathcal{V}$. We denote $f:\mathcal{V}^{L}\mapsto\mathbb{R}^{d}$ as the embedding function that maps sentences to the latent space, where $L$ is the maximum allowed words/tokens and $d$ is the dimension of latent embeddings. We use $\oplus$ to denote the concatenation of sentences. Following the convention, we will use the inner product of the embeddings $f(q)^{{\mathrm{\scriptscriptstyle T}}}f(p),$ to measure the similarity between two documents $p$ and $q$. We will use the $\angle(a,b)$ sign to denote the angle between two vectors $a$ and $b$.
Attacker’s capability There are three components for the retrieval systems in RAG: (1) data corpus, (2) retriever, and (3) query set. For the data corpus, we assume that the attacker can inject new data entries into it, e.g., by creating a new Wikipedia page or by entering a row of a fake patient’s information into an existing medical database. Regarding the retriever, we assume that the attacker can query the retriever models and view the retrieved documents and their associated latent embeddings. However, the attacker can neither speculate nor modify the parameters of the retriever models. For the query set, we assume that the attacker has access to queries of interests. In addition, we also examine the scenario where the attacker lacks access to the exact queries but has access to their semantically equivalent counterparts, which makes the considered threat model even more practical.
We provide two examples to demonstrate the feasibility of the above mentioned cases. The first example involves cases where a RAG system retrieves information from a public, large-scale corpus, such as Wikipedia [19, 15] and PubMed. In such instances, anyone, including the attacker, can edit the Wiki page and potentially exploit the threat model discussed above. The second example concerns the multi-modal large language model agents scenarios [17], where an agent aiming to address a question seeks assistance from another agent by retrieving relevant information from the latter’s local data corpus. In this scenario, a malicious agent has the capability to inject a few pieces of poisoned documents into their local data corpus in order to poison other agents’ decisions if using their poisoned information.
Attacker’s Goal Given an targeted document $T\in\mathcal{V}^{S}$ ($S\leq L$) and a set of queries $Q=\{q_{1},q_{2},\ldots,q_{n}\}$ with $q_{i}\in\mathcal{V}^{S}$, the attacker’s goal is to ensure that $T$ will consistently be retrieved with high ranking after injecting it into the data corpus, corresponding to attacker-specified queries $Q$. These types of goals are commonly observed in adversarial ranking/recommendations [18], where an attacker aims to improve the ranking of their targeted information.
3 Exposing the Safety Risks: A Case Study on Medical Retrieval Systems
@@ 3.1 Setups
Medical Corpus Following [14], we select a total of three medical-related corpora: (1) Textbook [20] ($\sim 126$K documents), containing medical-specific knowledge, (2) StatPearls ($\sim 301$K documents), utilized for clinical decision support, and (3) PubMed ($\sim 2$M documents), which consists of biomedical abstracts. Due to limited computation resources, the PubMed we used is a random subset of the total $23$M documents. Examples of each corpus and details about them are included in the appendix.
Targeted Information We consider a total of five types of targeted documents: synthetic personal identifiable information (PII), synthetic medical diagnose information, and adversarial passages generated (from [15]) for answering questions from for MA-MARCO [37], NQ [38], and HotpotQA [39], respectively. We use GPT-3 to evaluate their semantic closeness and conclude that they are semantically distant from each other. Therefore, we believe this setup encompasses a wide range of topics, strengthening the validity of our results. Examples are provided in the appendix.
Retriever We select three representative dense retrievers: (1) a general-domain semantic retriever: Contriever [29], (2) a scientific-domain retriever: SPECTER [30], and (3) a biomedical-domain retriever: MedCPT [31]. Details regarding these retrievers are included in the appendix.
@@ 3.2 Results
We report the success rates over a total of 225 combinations of query, corpus, retriever, and targeted information in Table 1 below. For each category of targeted information, we generated three different documents, calculated their success rates, and reported the mean value. The standard deviations are less than $0.07$. The interpretation of results in each cell adheres to the same following rule. We use the top-left cell as an example: it represents a success rate of $0.78$ achieved using the Corpus: Textbook, Retriever: MedCPT, Query set: MMLU-Med, with PPI as the targeted information.
@@
How robust is the attack against paraphrasing? There is one potential limitation regarding the above proposed attack through concatenation. For certain use cases in practice, the attacker may not know the exact queries. As a result, there could be no precise match between the queries used for retrieval and the queries used for creating poisoned documents. In the following, we demonstrate that the proposed attack is robust under such a mismatch. We maintain the same setup except that the queries are now rephrased by GPT-4 [2] to reflect the scenarios. We summarize the results in Table 2 below. We observed that the proposed attack remains effective under query paraphrasing, achieving a top-$2$ retrieval success rate of $0.8$ over most cases. These findings highlight even more pronounced security risks for retrieval in medical Q&A, as the attacker now only needs to know queries up to rephrasing in order to launch targeted attacks.
Table 1: Top 2 retrieval success rates over 3 corpora, 3 retrievers, 5 query sets, and 5 sets of targeted information. The interpretation of results in each cell adheres the following rule. We use the top-left cell as an example: it represents a success rate of $0.78$ achieved using the Corpus: Textbook, Retriever: MedCPT, Query set: MMLU-Med, with PPI as the targeted information.
@@
Corpus Retriever Query Target Information
PPI MS-MARCO NQ HotpotQA Diagnostic
Textbook MedCPT MMLU $0.76$ $0.72$ $0.61$ $0.52$ $0.64$
MedQA $0.98$ $0.97$ $0.98$ $0.97$ $0.96$
PubMedQA $0.94$ $0.96$ $0.92$ $0.95$ $0.91$
Specter MMLU $0.79$ $0.75$ $0.54$ $0.84$ $0.72$
MedQA $0.97$ $0.86$ $0.87$ $0.97$ $0.91$
PubMedQA $0.94$ $0.80$ $0.85$ $0.98$ $0.82$
Contriever MMLU $0.57$ $0.75$ $0.64$ $0.85$ $0.70$
MedQA $0.90$ $0.91$ $0.94$ $0.96$ $0.90$
PubMedQA $0.97$ $0.98$ $0.98$ $0.98$ $0.92$

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has been empirically shown to enhance the performance of large language models (LLMs) in knowledge-intensive domains such as healthcare, finance, and legal contexts. Given a query, RAG retrieves relevant documents from a corpus and integrates them into the LLMs’ generation process. In this study, we investigate the adversarial robustness of RAG, focusing specifically on examining the retrieval system. First, across 225 different setup combinations of corpus, retriever, query, and targeted information, we show that retrieval systems are vulnerable to universal poisoning attacks in medical Q&A. In such attacks, adversaries generate poisoned documents containing a broad spectrum of targeted information, such as personally identifiable information. When these poisoned documents are inserted into a corpus, they can be accurately retrieved by any users, as long as attacker-specified queries are used. To understand this vulnerability, we discovered that the deviation from the query’s embedding to that of the poisoned document tends to follow a pattern in which the high similarity between the poisoned document and the query is retained, thereby enabling precise retrieval. Based on these findings, we develop a new detection-based defense. Through extensive experiments spanning various Q&A domains, we observed that our proposed method consistently achieves excellent detection rates in nearly all cases.
@@ 1 Introduction
Large Language Models (LLMs) have achieved exceptional performance across a wide range of benchmark tasks spanning multiple domains [1, 2, 3]. However, they also suffer from several undesirable behaviors. For example, LLMs can generate responses that seem reasonable but are not factually correct, a phenomenon known as hallucination [4]. Additionally, due to data privacy regulations such as GDPR [5], direct training on specific data domains may be restricted. This can result in disparities between their acquired internal knowledge and the real-world challenges they encounter. It can be particularly concerning in domains require extensive knowledge, such as healthcare [6, 7], finance [8] and legal question-answering [9]. These challenges have sparked interest in more principled methods for decoding and alignment [10], as well as techniques that inject external knowledge to bridge domain gaps, such as Retrieval-Augmented Generation (RAG) [11, 12, 13, 14].
The RAG approach typically involves two steps: retrieval and augmentation. Upon receiving an input query, RAG retrieves the top $K$ relevant data from an external data corpus. It then integrates this retrieved information with its internal knowledge to make final predictions. Empirical evidence suggests that LLMs employing the RAG scheme significantly outperform their non-retrieval-based counterparts in knowledge-intensive domains like finance and medicine [13, 14]. For instance, the authors of [14] developed a state-of-the-art benchmark for the use of RAG in the medical domain. The authors observed an increase in prediction accuracy of up to 18% with RAG compared to non-retrieval and chain-of-thoughts versions across large-scale healthcare tasks, utilizing 41 different combinations of medical data corpora, retrievers, and LLMs.
@@ 1.1 Main Contributions
Revealing the safety risks for retrieval systems: case studies for medical Q&A and legal Q&A. We demonstrate that dense retrieval systems are vulnerable to what we term ‘universal poisoning attacks’ in medical Q&A across 225 use-case combinations of corpus, retriever, query, and targeted information. Similar observations are made for legal Q&A. As shown in Figure 1 below, in these attacks, adversaries can append nearly every sort of information, such as personally identifiable information (PII) and adversarial treatment recommendations, to a set of attacker-specified queries. Once these poisoned documents are injected into a large-scale corpus, such as Wikipedia and PubMed, they can be accurately retrieved, often with high rankings, e.g., top $1$, using attacker-specified queries. Depending on attackers’ goals, these documents will lead to safety risks such as (1) leakage of PII, (2) adversarial recommendations for treatments, and (3) jailbreaking the LLM during the inference stage once used as context.
@@
Figure 2: Defense methods against universal poisoning attacks. Clean documents (collected from the Textbook [20]) and poisoned documents are indistinguishable using popular defenses, such as the $\ell_{2}$-norm of their embeddings. However, under our proposed distance measurements, a clear separation between clean and poisoned documents is observed (right-most figure).
@@ 1.2 Related Work
Data poisoning and backdoor attacks. Our investigation is closely related to recent advances in understanding and defending against backdoor attacks in both training and inference stages. The adaptability hypothesis [21] and statistical perspectives on poisoning [22] offer complementary insights into how learned representations can be manipulated under adversarial influence. Furthermore, unified detection frameworks [23] propose practical strategies for identifying such attacks at inference time, although these methods have not yet been extensively explored in retrieval-augmented generation pipelines.
@@
Retrieval-Augmented Generation (RAG) RAG, popularized by [12], is a widely adopted approach that integrates retrieval-based mechanisms into generative models to improve performance across various language tasks [24]. The basic flow of RAG follows a ‘retrieve and read’ fashion [12, 25, 26, 13, 27, 11, 28, 29, 30]. In this fashion, given an input query, one first retrieves relevant data from the external data corpus and then employs generative models as ‘readers’, for example, by directly appending the retrieved documents to the queries as context and then feeding them as a whole to LLMs for making predictions. Approaches for enhancing the efficient use of retrieved documents to improve the LLMs’ reading capability include using chunked cross-attention during generation [13], prompt-tuning [26].
@@
Dense retrieval systems There are two main categories of retrievers: sparse retrieval systems, which match through lexical patterns (e.g., BM25) [31, 32], and deep neural network-based dense retrievers, which match through semantic meanings [33, 34, 35]. Due to the near-exact token-level matching pattern, sparse retrievers’ performance has been shown to be worse than that of dense retrievers in several domains [36], such as healthcare [14]. There are two popular approaches for training dense retrievers: supervised [37, 38] and self-supervised [33, 34, 35]. In the supervised regime, given a paired query and document, the goal is to maximize the similarity, i.e., inner product, between their embeddings. Motivated by advances in unsupervised learning, recent methods have started to apply contrastive learning [39] for training, where they observed improved performance across multiple benchmarks. In our work, we additionally consider a retriever specifically trained on medical datasets: MedCPT [35].
@@
Adversarial attacks against RAG/retrieval systems Current approaches for attacking RAG [15, 16] mostly involve poisoning the corpus to deceive the retrieval system [18, 40] into retrieving the poisoned documents for adversarial purposes. The recent work of PoisonedRAG aims to backdoor the RAG by tricking the LLM into generating attacker-specified answers based on the retrieved attacker-crafted documents [15]. The authors develop both black-box and white-box attacks to achieve this goal. In black-box methods, they directly append the context for answering the question to the query and inject it into the corpus, similar to our method in implementation. However, their approaches differ from ours in terms of goals, insights, and application scenarios. First, our objective is to investigate and comprehend the robustness of retrieval systems employed in RAG. We achieve this by injecting various types of information, encompassing both irrelevant and relevant content, into the corpus and evaluating the ease or difficulty of retrieval. Their goal, however, is to inject only query-relevant context to deceive the LLM’s generation process upon retrieval. Second, in terms of insights, we provide explanations on the difficulty/easiness of the retrieval of different kinds of information, which is not covered in their work. Third, we focus on the application domain of healthcare, while they focus on the general Q&A setting.
@@ 2 Preliminary and Threat Model
Notations Denote the set of vocabulary to be considered as $\mathcal{V}$. We denote $f:\mathcal{V}^{L}\mapsto\mathbb{R}^{d}$ as the embedding function that maps sentences to the latent space, where $L$ is the maximum allowed words/tokens and $d$ is the dimension of latent embeddings. We use $\oplus$ to denote the concatenation of sentences. Following the convention, we will use the inner product of the embeddings $f(q)^{{\mathrm{\scriptscriptstyle T}}}f(p),$ to measure the similarity between two documents $p$ and $q$. We use the $\angle(a,b)$ sign to denote the angle between vectors $a$ and $b$.
Attacker’s capability There are three components for the retrieval systems in RAG: (1) data corpus, (2) retriever, and (3) query set. For the data corpus, we assume that the attacker can inject new data entries into it, e.g., by creating a new Wikipedia page or by entering a row of a fake patient’s information into an existing medical database. Regarding the retriever, we assume that the attacker can query the retriever models and view the retrieved documents and their associated latent embeddings. However, the attacker can neither speculate nor modify the parameters of the retriever models. For the query set, we assume that the attacker has access to queries of interests. (We provide a detailed elaboration on this point, along with real-world examples, in Remark 2 in Section 3). In addition, we also examine the scenario where the attacker lacks access to the exact queries but has access to their semantically equivalent counterparts (see Table 7 in Section 3), which makes the considered threat model even more practical.
Remark 1 (Assumption on accessing the RAG database).
Accessing the database in RAG for medical scenarios is reasonable since publicly accessible databases have been frequently used in building RAG for medical use. For instance, the well-known PubMed, a comprehensive collection of biomedical literature citations and abstracts, has been frequently used for building RAG for medical use by researchers from NIH and Mayo Clinic [14, 41]. In particular, the work [14] demonstrated that by using PubMed as the knowledge database, the accuracy performance of RAG achieves better results than other medical texts, such as the Textbook [20]. PubMed is publicly accessible, which means an attacker could potentially inject adversarial information into it. This confirms the feasibility of our threat model.
Attacker’s Goal Given an targeted document $T\in\mathcal{V}^{S}$ ($S\leq L$) and a set of queries $Q=\{q_{1},q_{2},\ldots,q_{n}\}$ with $q_{i}\in\mathcal{V}^{S}$, the attacker’s goal is to ensure that $T$ will consistently be retrieved with high ranking after injecting it into the data corpus, corresponding to attacker-specified queries $Q$. These types of goals are commonly observed in adversarial recommendations [18], where an attacker aims to improve the ranking of their targeted information.
3 Exploring Safety Risks: Insights from Medical Retrieval Systems
@@ 3.1 Setups
Medical Corpus Following [14], we select a total of three medical-related corpora: (1) Textbook [20] ($\sim 126$K documents), containing medical-specific knowledge, (2) StatPearls ($\sim 301$K documents), utilized for clinical decision support, and (3) PubMed ($\sim 2$M documents), which consists of biomedical abstracts. Due to limited computation resources, the PubMed we used is a random subset of the total $23$M documents. Examples of each corpus and more details are included in the appendix.
Legal Corpus For the legal Q&A, we follow [9], where Case-Based Reasoning (CBR) is used to structure retrieval within the RAG process of an LLM to enhance legal Q&A performance. We use the Australian Open Legal QA (ALQA) dataset [42], with an example shown in Table 10 in the appendix.
Targeted Information (I) For the medical corpus: We consider five types of targeted documents: synthetic personally identifiable information (PII), synthetic medical diagnosis information, and adversarial passages generated from [15] for answering questions from MS MARCO [43], Natural Questions (NQ)[44], and HotpotQA[45], respectively. We use GPT-3 to evaluate their semantic similarity and find that they are semantically distant from each other. This suggests that our setup covers a broad range of topics, thereby enhancing the validity of our results. Examples are provided in the appendix.
(II) For the legal corpus: In addition to the aforementioned synthetic PII, HotpotQA, and NQ, we also prompt GPT to generate two types of synthetic legal content.
Retriever We select three representative dense retrievers: (1) a general-domain semantic retriever: Contriever [34], (2) a scientific-domain retriever: SPECTER [33], and (3) a biomedical-domain retriever: MedCPT [35]. Details on these retrievers are in the appendix.
@@ 3.2 Results
For the medical application, we report the success rates over a total of 225 combinations of query, corpus, retriever, and targeted information in Table 1 below. For the legal Q&A case, we report the attack success rates in Table 2. For each category of targeted information, we generated three different documents, calculated their success rates, and reported the mean value. The standard deviations are less than $0.07$. The interpretation of results in each cell adheres to the same following rule. We use the top-left cell as an example: it represents a success rate of $0.78$ achieved using the Corpus: Textbook, Retriever: MedCPT, Query set: MMLU-Med, with PPI as the targeted information.
@@
How robust is the attack against paraphrasing? There is one potential limitation regarding the proposed attack through concatenation. In certain practical use cases, users or defenders who are aware of the proposed attack may simply paraphrase the queries to defend against it. As a result, there may be no precise match between the queries used for retrieval and those used by attackers to create poisoned documents. In the following, we demonstrate that the proposed attack is robust under such a mismatch. We maintain the same setup except that the queries are now rephrased by GPT-4 [2] to reflect the scenarios. We summarize the results in Table 7 below. We observed that the proposed attack remains effective under query paraphrasing, achieving a top-$2$ retrieval success rate of $0.8$ over most cases. These findings highlight even more pronounced security risks for retrieval in medical Q&A, as the attacker now only needs to know queries up to rephrasing in order to launch targeted attacks.
Remark 2 (Further discussion on the assumption on accessing the query set).
Our proposed attacks are a type of RAG attack that uses a pre-selected set of queries (called targeted queries) as triggers to retrieve adversarial documents and poison the LLM generation pipelines. Some existing examples of these types of attacks can be found in references [15, 40].
In this scenario, attackers already know the queries since they pre-select them and design the poisoned documents, allowing them to launch the attack. The practical concern is whether normal users, who are unaware of the risks, will use these attacker-selected targeted queries. For example, if the queries only contain special but non-informative tokens, they are unlikely to be used by normal users. Therefore, to effectively target a large number of normal users, attackers need to identify the queries that normal users are likely to use and then create poisoned documents based on those queries.
In the following, we explain and provide new empirical results to show that the previously mentioned requirement can potentially be met in medical domains. We paraphrase all the normal queries to serve as a defense for the case where the attack knows exact queries. We observed that there is only a slight decrease in terms of attack success rates (detailed results can be found in Table 7 in the appendix). In other words, the attacker does not need to know the exact query. Instead, knowing queries with similar meanings/structures is sufficient for launching successful attacks. This requirement is relatively straightforward to fulfill because real-world medical queries tend to follow very standard patterns, which can be easily exploited by attackers.
To further demonstrate this point, we conducted a new experiment as follows. We employed GPT-4 to generate 100 new synthetic queries that are potentially commonly asked by doctors, based on the MedMCQA dataset. We repeated the same attack procedure described in the paper and observed an average attack success rate of approximately 0.92 for Contriever as the embedding model using the PubMed corpus. This result indicates that it is very easy for attackers to create meaningful medical queries that can lead to high attack success rates, which also shows that the RAG for medical use is at significant safety and security risks.
Table 1: Top 2 retrieval success rates over 3 corpora, 3 retrievers, 5 query sets, and 5 sets of targeted information (for the medical application). The interpretation of results in each cell adheres the following rule. We use the top-left cell as an example: it represents a success rate of $0.78$ achieved using the Corpus: Textbook, Retriever: MedCPT, Query set: MMLU-Med, with PPI as the targeted information.
@@
Table 2: Attack success rates (Top 2 retrieval rates) on legal Q&A.
Target Information
Task Retriever PII HotPotQA NQ Syn Legal I Syn Legal II
Legal Q&A Contriever 0.82 0.88 0.84 0.88 0.81
Specter 0.78 0.81 0.89 0.90 0.92